# Leveraging AI to Accelerate Clinical Data Cleaning: A Comparative Study of AI-Assisted vs. Traditional Methods

### v1

@@ 3.5. Usability and User Experience
Post-study usability assessments revealed unanimous preference for the AI-assisted system across all evaluated dimensions. NASA-TLX scores indicated significantly reduced perceived workload with the automated approach, with participants reporting lower mental demand, temporal pressure, effort requirements, and frustration levels compared to traditional methods. System Usability Scale scores for the AI system were on average 86.3, indicating excellent usability, while traditional spreadsheet methods scored 56 on average, reflecting substantial user experience challenges.
Qualitative feedback consistently highlighted the AI system’s ability to prioritize findings, provide contextual clinical information, and reduce the cognitive burden associated with cross-referencing multiple data sources. Participants particularly valued the system’s capability to surface complex cross-domain discrepancies that would be difficult to identify through manual review. Several participants noted that the AI system helped them identify discrepancy types they had not previously recognized, suggesting potential educational benefits beyond immediate productivity gains.
@@ 4. Discussion
Our findings demonstrate that AI assistance can fundamentally transform clinical trial data cleaning from a rate-limiting bottleneck into an accelerator of drug development. The magnitude of improvement observed, over 6-fold increase in throughput coupled with doubling of accuracy, exceeds typical incremental advances in clinical trial operations and suggests a true paradigm shift in how data quality assurance can be conducted.
4.1. Transforming Clinical Trial Timelines Through Enhanced Data Review Efficiency
The demonstrated 6-fold improvement in data review throughput achieved through AI-assisted clinical data cleaning has profound implications for accelerating clinical trial timelines across multiple critical phases. Traditional data cleaning bottlenecks that routinely delay trial milestones by weeks or months could be substantially reduced. This acceleration is particularly crucial for rare disease trials where patient populations face limited therapeutic options and urgent medical needs that make time-to-market considerations a matter of life and death.
The efficiency gains observed in this study directly address the most time-sensitive aspects of clinical trial conduct. Database preparation for interim analyses, which currently requires 4-8 weeks of intensive data cleaning effort, could potentially be completed within 1-2 weeks using AI-assisted methods. This compression of data review timelines enables more frequent and timely interim analyses, allowing Data Safety Monitoring Boards to make critical go/no-go decisions with more current information and potentially identifying safety signals or efficacy trends earlier in the trial lifecycle. For adaptive trial designs that rely on interim data to modify study parameters, the ability to rapidly prepare clean datasets becomes essential for maintaining trial momentum and maximizing the value of adaptive design features.
4.2. Revolutionizing Database Lock Procedures and Regulatory Submission Timelines
Database lock represents one of the most critical bottlenecks in running a clinical trial, where comprehensive data cleaning must achieve near-perfect quality standards before final analysis and regulatory submission. The 46.2% improvement in accuracy demonstrated by AI-assisted review could reduce the iterative cycles of query generation and resolution that traditionally extend database lock timelines by weeks. Current industry practice typically requires 7 weeks between last patient visit and database lock (Harper et al., 2021). AI-assisted review could potentially reduce this timeline to 1-2 weeks, representing a 75% reduction in database lock duration.
This acceleration has cascading effects throughout the regulatory submission process. Earlier database lock enables earlier statistical analysis, regulatory document preparation, and submission filing, potentially advancing market approval timelines by 3-6 months. For breakthrough therapies addressing unmet medical needs, this timeline compression could translate to thousands of additional patients gaining access to life-saving treatments months earlier than would be possible with traditional data management approaches. The economic implications are equally significant, with reduced development timelines extending patent life and improving return on investment for pharmaceutical sponsors.
4.3. Enabling Dynamic and Adaptive Trial Designs
The efficiency and accuracy improvements demonstrated by AI-assisted data review create new possibilities for implementing dynamic trial designs that were previously impractical due to data management constraints. Real-time data monitoring and continuous data cleaning become feasible when AI systems can maintain high-quality datasets with minimal human intervention. This capability enables seamless implementation of adaptive randomization, dose escalation decisions, and futility assessments that require rapid access to clean, analysis-ready data.
Platform trials and master protocols, which evaluate multiple treatments or indications simultaneously, could particularly benefit from AI-assisted data review. These complex trial designs generate massive datasets with intricate cross-study dependencies that challenge traditional manual review methods. The AI system’s demonstrated ability to identify cross-domain discrepancies and maintain consistency across multiple data sources makes it ideally suited for managing the complex data relationships inherent in platform trial designs. This could accelerate the adoption of more efficient trial designs that evaluate multiple therapeutic approaches in parallel rather than sequential monotherapy studies.
4.4. Reallocating Human Expertise to Higher-Value Clinical Activities
The dramatic efficiency improvements achieved through AI-assisted data review create opportunities to reallocate scarce clinical expertise from routine data checking activities to higher-value clinical interpretation and decision-making tasks. Clinical data managers and medical monitors can redirect their time from manual discrepancy detection to clinical assessment of identified issues, protocol deviation investigation, and strategic data quality planning. This shift from mechanical data review to clinical reasoning maximizes the value of expensive clinical expertise while maintaining rigorous oversight of automated processes.
The educational benefits observed in this study, where participants identified previously unrecognized discrepancy types, suggest that AI systems can enhance rather than replace human clinical judgment. By systematically flagging complex patterns and providing contextual clinical information, AI-assisted review can help clinical professionals develop deeper expertise in data quality assessment while simultaneously improving review efficiency. This human-AI collaboration model ensures that the clinical knowledge and regulatory experience essential for trial success remain central to data management processes while leveraging automation to eliminate routine, time-consuming tasks.
@@ 5. Conclusion
The integration of AI into clinical trial data cleaning represents a transformative opportunity to address fundamental inefficiencies in drug development. Our controlled experimental study provides robust evidence that AI assistance can simultaneously improve the speed, accuracy, and consistency of clinical data review while reducing operational burden on clinical trial sites.
@@
These findings have immediate practical implications for the pharmaceutical industry, suggesting that AI adoption could accelerate drug development timelines by weeks to months while improving data quality and reducing costs. More broadly, our results demonstrate the potential for human-AI collaboration to enhance performance in complex, safety-critical domains requiring both efficiency and expert judgment.
@@
As the pharmaceutical industry grapples with increasing trial complexity and data volumes, AI-assisted approaches offer a scalable solution that enhances rather than replaces human expertise. The consistent benefits observed across reviewer experience levels suggest that AI can democratize access to high-quality data review capabilities, potentially enabling more efficient global clinical trials.

### v2

@@ 4.5. Usability and User Experience
Post-study usability assessments revealed unanimous preference for the AI-assisted system across all participants. The inverse NASA Task Load Index (NASA-TLX) demonstrated a dramatic reduction in cognitive workload, with Octozi achieving a mean score of 83.3 compared to 25.7 for traditional spreadsheet methods (higher scores indicate reduced burden). This 3.2-fold improvement in task load represents a fundamental shift from high cognitive demand to manageable workload, with participants reporting substantial improvements across all six NASA-TLX dimensions: mental demand, physical demand, temporal demand, performance concerns, effort, and frustration.
System Usability Scale (SUS) assessments further validated the superior user experience of the AI-assisted approach. Octozi achieved a mean SUS score of 88.3, placing it in the ”A+” category (scores above 84.1), while traditional spreadsheet methods scored 55.4, falling into the ”D” category (scores between 51.7-62.6) (Lewis and Sauro, 2018). This 32.9-point difference represents not merely an incremental improvement but a transformation from a ”D” to an ”A+” user experience (Fig. 3).
Qualitative feedback consistently highlighted three key advantages of the AI system. First, participants valued the intelligent prioritization of findings, which directed attention to the most clinically significant discrepancies rather than requiring sequential review of all data points. Second, the provision of contextual clinical information eliminated the need for constant reference switching between multiple data sources, reducing both time and cognitive load. Third, the system’s ability to surface complex cross-domain discrepancies revealed patterns that participants acknowledged would be ”nearly impossible to detect through manual review in the same time,” as noted by one senior medical monitor with 6 years of experience.
@@ 5. Discussion
Our findings demonstrate that AI assistance can fundamentally transform medical data review from a rate-limiting bottleneck into an accelerator of drug development. The magnitude of improvement observed over 6-fold increase in throughput coupled with doubling of accuracy, exceeds typical incremental advances in clinical trial operations and suggests a true paradigm shift in how data quality assurance can be conducted.
5.1. Implications for Clinical Trial Operations
The demonstrated improvements have immediate and far-reaching implications for clinical trial conduct. The 6-fold improvement in data review throughput could compress critical trial milestones that currently delay drug development by weeks or months. Database preparation for interim analyses, which traditionally requires 4-8 weeks of intensive effort, could potentially be completed within 1-2 weeks using AI-assisted methods. This acceleration is particularly crucial for adaptive trial designs that rely on timely interim data to modify study parameters, and for rare disease trials where patient populations face urgent medical needs.
The 46.2% improvement in accuracy, combined with the 15.48-fold reduction in false positive queries, addresses two fundamental challenges in clinical data management: maintaining high data quality standards while minimizing site burden. Current industry practice shows that medical reviewers generate queries with only 51% accuracy (Stokman et al., 2021), creating substantial administrative burden for clinical sites. Our results demonstrate that AI assistance can nearly eliminate unnecessary queries while maintaining high sensitivity for true discrepancies, potentially transforming the relationship between sponsors and investigative sites.
Figure 5. Relationship between reviewer experience and performance improvement with AI augmentation. The positive trend line (r=0.655) indicates that more experienced reviewers achieved greater absolute improvements when using AI assistance. This pattern suggests that AI augmentation serves as a force multiplier for clinical expertise rather than merely a substitute for experience.
5.2. Economic Impact Analysis
To quantify the economic implications of our findings, we conducted a comprehensive cost-benefit analysis for a representative Phase III oncology trial with 1,100 patients generating approximately 4.5 million data points over four years. Table 2 summarizes the primary areas of cost savings achieved through AI-assisted data review. The analysis reveals three primary areas of economic impact:
Table 2. Economic impact analysis of AI-assisted medical data cleaning for a representative Phase III oncology trial (1,100 patients, 4.5M data points, 4-year duration). Conservative estimates assume 50% realization of observed experimental improvements to account for real-world implementation factors.
Cost Category Traditional Cost AI-Assisted Cost Savings % Reduction Basis of Calculation
Medical Review
Review hours 2,100 hrs 700 hrs 1,400 hrs 66.7% 3x efficiency gain
Total cost (\$300/hr) \$630,000 \$210,000 \$420,000 66.7% (conservative vs. 6.03x observed)
Query Management
Total queries generated 5,154 2,881 2,273 44.1% 10x false positive reduction
False positive queries 2,526 253 2,273 90.0% (conservative vs. 15.48x observed)
Total cost (\$126.50/query) \$651,949 \$364,439 \$287,510 44.1% 56% reduction in queries
Database Lock
Timeline (days) 36.8 31.8 5.0 13.6% Conservative estimate
Lost revenue (\$840K/day) \$30,912,000 \$26,712,000 \$4,200,000 13.6% (expert panel: 5-12 day range)
Operational cost (\$40K/day) \$1,472,000 \$1,272,000 \$200,000 13.6%
Subtotal \$32,384,000 \$27,984,000 \$4,400,000 13.6%
Total Direct Savings \$33,665,949 \$28,558,439 \$5,107,510 15.2%
Medical Review Efficiency: Traditional medical review requires approximately 2,100 hours across three reviewers, costing \$630,000. Applying a conservative 3-fold efficiency improvement (half of our observed 6.03-fold gain), total review time reduces to 700 hours, yielding \$420,000 in savings.
Query Management Optimization: With medical reviewers generating and managing 5,154 queries at \$126.50 each, traditional approaches cost \$651,949. Our demonstrated 15.48-fold reduction in false positives (conservatively adjusted to 10-fold) reduces total queries by 56%, saving \$287,510 while maintaining high sensitivity for true discrepancies.
Database Lock Acceleration: The most substantial savings derive from accelerated database lock. A conservative 5-day reduction from the typical 36.8-day timeline saves \$4.4 million through earlier market entry (\$840,000/day in opportunity cost) and reduced operational expenses (\$40,000/day).
The cumulative direct savings of \$5.1 million represent only quantifiable operational improvements. Additional unquantified benefits include improved regulatory submission quality, earlier safety signal detection, reduced protocol deviations, and enhanced site relationships through reduced query burden. A detailed economic analysis is provided in Appendix A.
5.3. Human-AI Collaboration Model
Our results provide compelling evidence for a collaborative model where AI augments rather than replaces human expertise. The positive correlation between reviewer experience and AI-enabled performance gains (r=0.655, Fig. 5) suggests that AI serves as a force multiplier for clinical expertise. Experienced reviewers achieved the greatest absolute improvements, gaining up to 28 additional correct predictions compared to 12 for novice reviewers, indicating that domain knowledge remains crucial for maximizing AI benefits.
This collaborative model addresses key concerns about AI adoption in regulated environments. By maintaining human oversight while automating routine tasks, the system preserves the clinical judgment essential for complex medical decisions while eliminating the mechanical aspects of data review. The educational benefits observed where participants identified previously unrecognized discrepancy types further suggest that AI systems can enhance professional development while improving operational efficiency.
@@ 6. Conclusion
The integration of AI into clinical trial data cleaning represents a transformative opportunity to address fundamental inefficiencies in drug development. Our controlled experimental study provides robust evidence that AI assistance can simultaneously improve the speed, accuracy, and consistency of medical data review while reducing operational burden on clinical trial sites. The observed 6.03-fold improvement in throughput, coupled with a doubling of accuracy and 15.48-fold reduction in false positive queries, demonstrates the potential for AI to accelerate drug development timelines by weeks to months while maintaining or improving data quality standards.
@@
These findings offer immediate practical implications for the pharmaceutical industry, with potential cost savings exceeding \$5 million per Phase III trial through improved efficiency, reduced query burden, and accelerated database lock. More broadly, our results demonstrate the potential for human-AI collaboration to enhance performance in complex, safety-critical domains requiring both efficiency and expert judgment.
@@
As the pharmaceutical industry grapples with increasing trial complexity and data volumes, AI-assisted approaches offer a scalable solution that enhances rather than replaces human expertise. The consistent benefits observed across reviewer experience levels suggest that AI can democratize access to high-quality data review capabilities, potentially enabling more efficient global clinical trials and ultimately accelerating the delivery of life-saving therapies to patients in need.
Appendix A Detailed Economic Impact Analysis
A.1. Methodology and Assumptions
This economic analysis evaluates the financial implications of implementing AI-assisted medical data cleaning in a representative Phase III oncology trial. Our model is based on:
  • Trial parameters: 1,100 patients, 4.5 million data points, 4-year duration
  • Industry benchmarks from published literature and proprietary databases
  • Expert panel consisting of three senior medical monitors (average 20+ years experience)
  • Conservative adjustment factors applied to experimental results
A.2. Detailed Cost Breakdown
A.2.1. Medical Review Labor Costs
Traditional medical review requirements were calculated based on insight from the expert panel:
  • Baseline workload: 150-200 hours per reviewer per year (used 175)
  • Team size: 3 medical reviewers (standard for oncology Phase III)
  • Total hours: 2,100 hours over 4 years
  • Hourly rate: \$300 (rate for an experienced senior medical monitor at a sponsor)
  • Total traditional cost: \$630,000
AI-assisted projections assume a conservative 3-fold efficiency gain (50% of the 6.03-fold improvement observed experimentally), accounting for:
  • Data cleaning and review is typically done in batches
  • Rounding up data review hours over the course of trial
  • Estimated reviewer fatigue for long session over 1 hours
A.2.2. Query Management Economics
Query generation and resolution represents a significant cost center:
  • Total data points: 4,341,900 (based on 4100 fields per patient and 1100 patients in this study)
  • Manual query rate: 20% of queries are estimated to be manual as opposed to automated edit checks (Lewis and Sauro, 2018)
  • Industry query rate: 7.3% of data points for oncology study (Stokman et al., 2021)
  • Medical reviewer queries: 8.13% of total queries (Stokman et al., 2021)
  • Cost per query: \$126.50 (average of \$28-\$225 range) (Stokman et al., 2021)
Query cost components include:
  • Initial query generation and documentation
  • Clinical team review and approval
  • Site communication and follow-up
  • Response review and possible escalation
The 15.48-fold reduction in false positives observed experimentally was conservatively adjusted to 10-fold for real-world projections, considering:
  • Medical reviewers aiming to clarify information instead of disputing data
  • Estimated efficiency losses
The savings related to query management for medical reviewers considers:
  • % of no-change queries 51% of queries created by medical reviewers do not result in changes to the dataset (Stokman et al., 2021)
  • Number of false queries 2,525 (5,1540.515,154*0.51)
  • Reduction in false query rate 10-fold decrease (conservative estimate)
  • Remaining false queries 253 ($2,525/10$)
A.2.3. Database Lock Value Analysis
Database lock acceleration provides the highest economic value:
  • Industry average DBL: 36.8 days (range: 2-120 days) (Harper et al., 2021)
  • Conservative reduction: 5 days (expert panel estimated 5-12 days)
  • Daily opportunity cost: \$840,000 (median loss of revenue from delayed trial cost for oncology trial (Smith et al., 2024))
  • Daily operational cost: \$40,000 (clinical team costs, manufacturing, sites, etc. (Smith et al., 2024))
A.3. Indirect Benefits Not Quantified
Several additional benefits were not included in the financial analysis:
  • Improved safety signal detection: Earlier identification of adverse events could prevent costly late-stage failures (average Phase III failure costs \$150M)
  • Enhanced regulatory submission quality: Reduced likelihood of Complete Response Letters (CRL), each potentially costing \$50-100M in delays
  • Site relationship improvements: Reduced query burden may improve site retention and recruitment for future trials
  • Competitive advantage: Faster trial completion enables earlier market entry and extended market exclusivity
  • Portfolio effects: Efficiency gains enable reallocation of resources to additional pipeline programs
Total implementation investment of approximately \$400,000 is recovered within the first month of database lock acceleration.
Appendix B Enrollment Questionnaire
Table 3 describes the question asked to interested participants in the study and the type of response for that question. These questions help ascertain if a potential participant was eligible for the study, contact information, and information to confirm their background.
Table 3. Screening questions asked to respondents.
Question Field Type
First and last name Text
Email Text
LinkedIn Text
Resume / CV File
How would you describe your current or most recent role? (Select all that apply) Checkboxes
Does your current or most recent role involve medical cleaning or reviewing of clinical trial data? Yes/No
How many years have you worked in roles involving data cleaning or clinical data review? <2, 2-5, 5+ years
Roughly how many clinical trials have you been involved in where you were responsible for medical data cleaning? <2, 2-5, 5+ years
Have you had experience with medical data cleaning/review for oncology studies? Yes/No
Which tools have you used for data cleaning or review? (Select all that apply) Checkboxes
Why are you interested in participating in this study? Text
Appendix C Octozi Platform
Figures 6 and 7 are screenshots of the Octozi platform that a user would see after selecting a data entry (one case report entry) in a previous menu. What is not presented in these figures are sorting algorithms that present the user with critical entries that contain potential discrepancies first which lower priority and likely clean data last.
Figure 6. A screenshot of the Octozi platform for an adverse event. The page presents the user with a short narrative describing the context for the adverse event of anemia, discrepancies found by the platform, and all of the patient information as figures in ”Patient Profile” or in tabular form in ”Cross-Database Entries”.
Figure 7. A screenshot of the Octozi platform for a discrepancy check that has found a potential inconsistency in the data. The context engine has pulled relevant adverse event and laboratory case report forms with limited column values for the adverse event in question. Not shown are the CTCAE reference ranges and text explanations for anemia. The output of the AI-assistant is provided at the top of the reference information to direct the user to its conclusion.